the length of a rectangle is 8 inches more than the width. The perimeter is 36 inches. What are the length and width of the rectangle
Answer by Cromlix(4381)

Width = x
Length = x + 8
Perimeter = 2*length + 2*width
36ins = 2(x + 8) + 2(x)
36ins = 2x + 16 + 2x
36 - 16 = 4x
4x = 20
x = 5
Width = 5ins
Length = 13ins
Hope this helps.
:-)